1. How many ways can you make change for $50 bill using $5, $10 and $20 bills?
2. A compass and ruler together cost $4. The compass costs 90cents more than the ruler. How much does the compass cost?

Answer by checkley77(12844)

I GET 40 COMBINATIONS OF CHANGE.
20--10--5 (these are the bills)
-----------------
2 (these are the quantities of the various bills)
1
----4
----3
----2
----1
--------9
--------8
--------7
--------6
--------5
--------4
--------3
--------2
--------1
2-------1
1---2
1---2---1
1-------5
1-------4
1-------3
1-------2
1-------1
----4
----4---1
----3---3
----3---2
----3---1
----2---5
----2---4
----2---3
----2---2
----2---1
----1---7
----1---6
----1---5
----1---4
----1---3
----1---2
----1---1
--------------------------
r+c=4.00
c=r+.90
r+(r+.90)=4.00
2r+.90=4.00
2r=4.00-.90
2r=3.10
r=3.10/2
r=1.55 ANS.
c=1.55+.90
C=2.45 ANS.
PROOF:
1.55+(1.55+.90)=4.00
1.55+2.45=4.00
4.00=4.00